# CLAPS: Posterior-Aware Conformal Intervals via Last-Layer Laplace

### v1

@@ Appendix B Posterior-Variance Diagnostics and Heteroscedasticity Signals
Table 4: LLLA variance decomposition by dataset and split. Columns report the mean epistemic component, the fractional epistemic share $r=\mathrm{epi}/(\sigma^{2}+\mathrm{epi})$ (mean and 10/90th percentiles), the fraction of samples with $r<1\%$, the posterior covariance trace $\mathrm{tr}(\Sigma)$, and the aleatoric variance $\sigma^{2}$.
@@
Airfoil Calib 0.007292 0.075838 0.049370 0.105913 0.000000 32.587912 0.088094
Test 0.007931 0.080710 0.047274 0.108744 0.000000 32.587912 0.088094
kin8nm Calib 0.001572 0.018776 0.012259 0.027264 0.037940 5.269668 0.081985
Test 0.001548 0.018492 0.011573 0.026858 0.035366 5.269668 0.081985
CASP Calib 0.002007 0.005018 0.001983 0.008518 0.928086 65.819530 0.396034
Test 0.002074 0.005168 0.002061 0.008725 0.926088 65.819530 0.396034
YearPrediction Calib 0.000257 0.000471 0.000159 0.000882 0.999526 56.635682 0.545434
Test 0.000262 0.000480 0.000158 0.000892 0.999437 56.635682 0.545434
Table 5: Heteroscedasticity signal: Spearman correlation between $|e|$ and $\sqrt{\hat{v}}$. Two panels report calibration and test splits side by side.
Dataset $\rho$ $p$ Airfoil -0.0640 0.4592 kin8nm 0.0850 0.0210 CASP 0.0111 0.4746 Year 0.2271 0.0000
Dataset $\rho$ $p$ Airfoil 0.0073 0.9290 kin8nm 0.0489 0.1616 CASP 0.0103 0.4841 Year 0.2298 0.0000
Figure 2: Subsample diagnostics shown as epoch-style curves (1–5, interpolated per dataset on a log-$n$ grid). Panels report Epistemic (mean), $\mathrm{tr}(\Sigma)$, and $\hat{\sigma}^{2}$ from left to right, with four datasets overlaid in each panel (legend shared at top). This view makes small-dataset trends comparable across datasets with differing raw subsample sizes.

### v2

@@ Appendix B Posterior-Variance Diagnostics and Heteroscedasticity Signals
Table 4: LLLA variance decomposition by dataset and split (mean $\pm$ standard deviation over five random seeds). Columns report the mean epistemic component, the fractional epistemic share $r=\mathrm{epi}/(\sigma^{2}+\mathrm{epi})$ (mean and 10/90th percentiles), the fraction of samples with $r<1\%$, the posterior covariance trace $\mathrm{tr}(\Sigma)$, and the aleatoric variance $\sigma^{2}$.
@@
Airfoil Calib $0075\pm 0.0003 0.0$$764\pm 0.0016 0.0$$0488\pm 0.0015$ $1080\pm 0.0030 0.0$$0000\pm 0.0000$ $33.0799\pm 3.3503$ $0.0896\pm 0.0032$
Test $0077\pm 0.0003 0.0$$783\pm 0.0032 0.0$$0489\pm 0.0017$ $1114\pm 0.0069 0.0$$0000\pm 0.0000$ $33.0799\pm 3.3503$ $0.0896\pm 0.0032$
kin8nm Calib $0016\pm 0.0001 0.0$$84\pm 0.0005 0.011$$18\pm 0.0004 0.0$$261\pm 0.0009 0.0$$0358\pm 0.0075$ $6.5324\pm 1.7552$ $0.0854\pm 0.0044$
Test $0016\pm 0.0001 0.0$$82\pm 0.0003 0.011$$16\pm 0.0001 0.0$$262\pm 0.0007 0.0$$0380\pm 0.0038$ $6.5324\pm 1.7552$ $0.0854\pm 0.0044$
CASP Calib $0020\pm 0.0000 0.0$$051\pm 0.0002 0.0$$019\pm 0.0001 0.0$$0089\pm 0.0003$ $0.9215\pm 0.0065$ $66.2770\pm 2.3513$ $0.3909\pm 0.0079$
Test $0020\pm 0.0000 0.0$$051\pm 0.0001 0.0$$019\pm 0.0001 0.0$$0089\pm 0.0002$ $0.9211\pm 0.0061$ $66.2770\pm 2.3513$ $0.3909\pm 0.0079$
YearPrediction Calib $003\pm 0.0000 0.000$$5\pm 0.0000 0.000$$2\pm 0.0000 0.000$$009\pm 0.0000$ $0.9995\pm 0.0001$ $59.0817\pm 8.2215$ $0.5527\pm 0.0115$
Test $003\pm 0.0000 0.000$$5\pm 0.0000 0.000$$2\pm 0.0000 0.000$$009\pm 0.0000$ $0.9995\pm 0.0001$ $59.0817\pm 8.2215$ $0.5527\pm 0.0115$
Table 5: Heteroscedasticity signal: Spearman correlation between $|e|$ and $\sqrt{\hat{v}}$. Two panels report calibration and test splits side by side. Entries show mean $\pm$ standard deviation over five random seeds for the correlation coefficient $\rho$ and its $p$-value.
Dataset $\rho$ $p$ Airfoil $0.0057\pm 0.0690$ $0.5685\pm 0.2552$ kin8nm $0890\pm 0.0461 0.0$$0971\pm 0.1236$ CASP $0.0137\pm 0.0077$ $0.4255\pm 0.2509$ Year $2466\pm 0.0161 0.0$$0000\pm 0.0000$
Dataset $\rho$ $p$ Airfoil $0.0029\pm 0.0898$ $0.4651\pm 0.3319$ kin8nm $1043\pm 0.0557 0.0$$0418\pm 0.0681$ CASP $0.0071\pm 0.0195$ $0.3253\pm 0.3082$ Year $2442\pm 0.0139 0.0$$0000\pm 0.0000$
Figure 2: Subsample diagnostics for LLLA shown as epoch-style curves over five subsample sizes (1–5, spaced on a log-$n$ grid). Each panel reports one quantity—Epistemic (mean), $\mathrm{tr}(\Sigma)$, and $\hat{\sigma}^{2}$ from left to right—with all four datasets overlaid and a shared legend at the top. Curves are plotted for a single representative random seed and are intended to illustrate qualitative posterior behavior across datasets with different absolute sample sizes.